OpenStudy (anonymous):

sqrt x+3-3=x Solve for x???? ( confused again :/)

OpenStudy (phoenixfire):

\[\sqrt{x+3} - 3 = x\] ?

OpenStudy (anonymous):

yeah thats the question :/

OpenStudy (anonymous):

Add 3 first both the sides and tell me what did you get ??

OpenStudy (anonymous):

sqrt x+3= x+3

OpenStudy (phoenixfire):

Now square both sides.

OpenStudy (anonymous):

\[(\sqrt{x+3})^2 = (x + 3)^2\]

OpenStudy (phoenixfire):

sqrt and the square cancel. expand the brackets on the right.

OpenStudy (anonymous):

No need to square I guess..

OpenStudy (anonymous):

Britney do you have answers with you ??

OpenStudy (anonymous):

my options are: x=-2 x=-3 x=-2 and -3 None of the above

OpenStudy (anonymous):

o.O

OpenStudy (phoenixfire):

\[\sqrt{x+3}-3 = x\] \[\sqrt{x+3}=x+3\] \[x+3=(x+3)^2\] \[x+3=x^2+6x+9\] \[0 = x^2+6x+9-x-3\] \[0=x^2+5x+6\] apply quadratic formula \[{-b \pm \sqrt{b^2-4ac}} \over 2a\] a=1, b=5, c=6

OpenStudy (anonymous):

I got